I would like some help please with the following problem:
A mixture contains 25% syrup and 75% water. If the amount of water is tripled, what percentage of the new mixture is syrup?

Answer by ankor@dixie-net.com(12701)

A mixture contains 25% syrup and 75% water. If the amount of water is tripled, what percentage of the new mixture is syrup?
:
Let's say the original amt = 12 gal
That makes the amt of water = .75(12) = 9 gals, 3 times the water = 27 gals
We still have .25(12) = 3 gal of syrup, + 27 gal of water; A total of 30 gal
:
Write a per cent syrup equation:
.25(12) = x(30)
3 = 30x
x = 3/30 = .1
:
You have 10% syrup if you triple the amt of water.